OpenStudy (anonymous):

@countonme123 At the beginning of the day, the sum of the stocks in the Dow Jones was $10,245.31. At the end of the day, the sum of the stocks in the Dow Jones was $10,165.79. What was the rate of change of the Dow Jones on this particular day? A. –1.2% B. 0.8% C. –0.8% D. –1.0% @dss08a @Euler271 @Tiroc @Zilther

OpenStudy (anonymous):

just by looking at the answers i would think is -.8

OpenStudy (anonymous):

 find the measure of the angle

OpenStudy (anonymous):

the sum of the angles in a rectangle is 360 so 360-109-100-90 = ?

OpenStudy (anonymous):

61 degrees ?

OpenStudy (anonymous):

@divagirl421 10,245.37 - 10,165.79 = 79.58 this is the amount it decreased and 79.58/10,245.37 = .0078 and .0078 = .008 = .8% and since the number decreased then it has to be negative so the final answer is -.8%

OpenStudy (anonymous):

@Bbambino1976 yes

OpenStudy (anonymous):

thanks I knew that lol

OpenStudy (anonymous):

 how do I find m<t? help!!

OpenStudy (anonymous):

type the question in a different conversation and i will help you

OpenStudy (anonymous):

anyways in a parallelogram opposite angles are equal so 10x+5 = 49x -2 solve for x and you get that x = .18 so those angles are 10(.18)+5 =6.8 and like i said a shape with four angles has 360 degrees and so 360 - 6.8*2 = 346.4 and this divided by two gives you your angle so 346.2/2 = 173.1 and that is your answer

OpenStudy (anonymous):

lol this is econ not geo the answer is -1.6 -apex

OpenStudy (whpalmer4):

To find the percentage change going from A to B, compute: \[\frac{B-A}{A}*100\%\] So here we have \[A = 10,245.31\]\[B=10,165.79\] \[\frac{B-A}{A}*100\%=\frac{10165.79-10245.31}{10245.31}*100\%=\frac{-79.52}{10245.31}*100\%=\] Hint for @sjohnson4772: when your answer is not one of the answer choices in a multiple-choice problem, recheck your work!